# PBEBench: A Multi-Step Programming by Examples Reasoning Benchmark inspired by Historical Linguistics

### v3

@@ 5.2 PBEBench Performance
We evaluate the best open and closed-source models, gpt-oss-120b and GPT-5, on PBEBench, which includes cascades of length 2–20 and harder snapshots of length 25 and 30. On cascades of length 2–20, gpt-oss-120b achieves Pass@1 of 0.67, average Edit_Sim of 0.95, and Valid_Rate of 0.96. Per cascade Pass@1 are values added in (Table 11). GPT-5, being stronger, achieves Pass@1 of 0.94 on cascades of length 10 with just 1 sample (versus Pass@1 of 0.84 with 32 for gpt-oss-120b) and 0.44 on cascades of length 20 with 4 samples (versus Pass@1 of 0.05 with 32 for gpt-oss-120b). Since GPT-5 still performs meaningfully at length 20, unlike gpt-oss-120b which collapses to 5%, we evaluate it on harder cascades of length 20, 25, and 30, where it averages Pass@1 of 0.175 (Table 13). Pass@1 trends are shown in Fig. 2(a) (gpt-oss-120b, 2–20) and Fig. 2(b) (GPT-5, 20–30). Both models follow the scaling strategies in Section 3.2: gpt-oss-120b uses sampling budget 32 with max sequence length 16384, while GPT-5 uses sampling budget 4 (to mitigate linear cost scaling), a completion limit of 65536 tokens (including reasoning and output tokens), and moderate reasoning effort to make sure the reasoning terminates. Performance declines nearly monotonically with increasing cascade length: gpt-oss-120b drops to about 50% and 5% at lengths 15 and 20, while GPT-5 drops to 14% and 5% at 25 and 30, highlighting reasoning limits in both. Logistic regression on gpt-oss-120b shows cascade length is a strong negative predictor of Pass@1, with feeding, counter-feeding, and counter-bleeding also significant predictors of failure and bleeding associated with success (Table 10).
5.3 Ablations
We evaluate gpt-oss-20b and gpt-oss-120b on the PBEBench-Lite-MoreEg snapshot (Section 1) and show the results in Table 12. The results show similar to slightly lower performance, which is counterintuitive because historical linguists anecdotally perform better with more examples. We analyze the fraction of examples modified by each program for both PBEBench-Lite and PBEBench-Lite-MoreEg and find that it is 1.56/5 (31%) and 7.07/50 (14%) for each dataset, meaning that while there is greater absolute effect of each program in the latter dataset, the relative effect if lesser (ratio of changed to total examples). This might explain the difficulty for PBEBench-Lite-MoreEg. We also evaluate the effect of scaling the sampling budget for a constant max sequence length and vice versa in Fig 3(a) (Table 16) and Fig 3(b) (Table 18) respectively for gpt-oss-120b and Fig 4(a) (Table 14) and Fig 4(b) (Table 15) respectively for GPT-5 (we compute k-fold average for GPT-5 to account for variance, exact details of which are provided in E.7). All the plots show rapid initial gains followed by a regime of diminishing returns and eventual saturation.
(a) Pass@1 vs Sampling Budget (max seq len: 16384)
(b) Pass@1 vs Max Seq Len (sampling budget: 32)
(b) Pass@1 vs Max Completion Tokens (sampling budget: 1)
@@ 6 Discussion
After analyzing the results, we address the research questions from Section 1. RQ1: Which models are good at inductive reasoning?: Long chain-of-thought models consistently perform better across both open and closed-source categories on PBEBench-Lite, with gpt-oss-120b and GPT-5 being the best in each. RQ2: What makes a PBE instance hard?: Factorial and regression analysis (Appendix G.3G.4) show cascade length as the most influential factor, followed by certain BFCC relations (especially feeding and counter-feeding). Counterintuitively, more examples reduce performance, as the relative density of updates per program decreases with increasing examples (see Sec. 5.3). RQ3: Can the strongest LLMs solve problems requiring long programs (e.g., real-world forward reconstruction)?: Even the strongest LLMs struggle on PBEBench, despite large scaling budgets. Real world forward reconstruction problems are even harder, with higher cascade lengths and more examples. RQ4: What scaling strategies work best for solution search?: A minimum value for max sequence length is critical for reasoning models, and this requirement grows with model capability. Without it, performance drops to zero. GPT-5, being more capable, has a higher minimum requirement than gpt-oss-120b. Increasing reasoning_effort for GPT-5 further raises this requirement, and with a fixed max sequence length, performance declines due to more unterminated outputs (Fig. 18). Beyond this threshold, either scaling strategy (sampling budget and max sequence length) is roughly equivalent for gpt-oss-120b (e.g., 32768 max sequence length with sampling budget 32 $\approx$ sampling budget 64 at 16384). However, scaling sampling budget is slower if not parallelized and more redundant, as attempts cannot share steps, whereas a single long chain of thought can. A balanced use of both strategies is therefore necessary.
@@ 7 Conclusion and Future Work
PBEBench-Lite is already challenging for many models, especially non-reasoning ones, indicating that inductive reasoning requires long chains of thought at test time. Experiments with the strongest open and closed-source models on PBEBench show that while they handle moderate complexity, performance steadily declines as complexity, mainly driven by ground-truth cascade length, increases. Even with intensive scaling strategies like larger sampling and token budgets, performance eventually saturates, highlighting the reasoning limits of current state-of-the-art models. Nonetheless, results are promising and surpass prior work (Naik et al., 2024; 2025b), pointing to a bright future for solving forward reconstruction problems with LLMs. Moreover, our automated, scalable data generation with controllable complexity could enable increasingly complex synthetic training data, potentially enhancing PBE and inductive reasoning in open-weight models via curriculum-style learning, an avenue we plan to pursue in future work.
8 Ethics Statement
Our work introduces a provably correct, fully automated, and scalable pipeline for generating multi-step Programming by Example (PBE) problems, enabling the evaluation of inductive reasoning in Large Language Models without human supervision. As no human subjects are involved, we do not anticipate ethical concerns. Moreover, since the dataset focuses solely on assessing abstract inductive reasoning in LLMs, it is unlikely to be misused for harmful purposes.
9 Reproducibility Statement
To facilitate reproduction of our results, we thoroughly document all components of our work. The BFCC relation types, their detectors (Algorithm 2), and proofs of correctness are presented in Appendix C. The rejection-sampling-based data generation process is both described (Section 3.1) and formally specified (Algorithm 1). Detailed statistics and parameters for all generated datasets appear in Section 4.1 and Appendix D.2. Our prompting and program extraction strategy, including scaling techniques, are reported in Section 3.2, while the prompt template is provided in Appendix E.2, and inference parameters for each model are given in Appendix F.2. For closed source models we use specific dated snapshots/checkpoints of models as recorded in Appendix E.6 to aid with reproducibility. To estimate the amount of variance between runs/experiments, we do a small controlled experiment for gpt-oss-120b (Table 17), and compute k-fold average for GPT-5 (Appendix E.7) The evaluation procedure and metrics are described in Section 4.3 and Section 3.2. Finally, we discuss limitations of our work in Appendix B.
Appendix A Appendix
@@
    Appendix B presents the limitations.
@@
    Appendix C provides the theory and proofs for the relation type classifiers used to control the distribution.
@@
    Appendix D details the benchmark, including licensing, statistics, and related information.
@@
    Appendix E gives additional methodological details, including prompts, model selection, licensing, run costs, algorithm pseudocode, snapshots for closed-source models, and the K-fold evaluation used to simulate GPT-5’s sampling budget.
@@
    Appendix F outlines experimental details such as computational environment, inference parameters, and strategies for CoT truncation/control with gpt-oss-120b.
@@
    Appendix G presents additional results, including data generation efficiency, factorial and logistic regression analyses of instance difficulty, confusion matrices comparing dataset ground truth and model predictions, and extended tables for scaling experiments, effects of additional examples, and performance breakdowns by cascade length.
Appendix B Limitations
Despite this paper’s significant contributions, such as a scalable and provably correct dynamic data generation pipeline for instances of controllable complexity and extensive benchmarking of state-of-the-art reasoning models under popular scaling strategies, several limitations remain:
  1. 1.
    Due to cost, we could not use very large sampling budgets with GPT-5, nor evaluate it on every cascade length, restricting experiments to the most challenging cases where performance drops to nearly zero.
  2. 2.
    Due to randomness in the sampling procedure, it cannot fully balance constraints such as relationship types and cascade lengths, especially for very long cascades, where suitable examples are increasingly rare, leading to high rejection rates and exhaustion of patience $\tau$. Future work could study the impact of targeted sampling, using algorithmic approaches to balance constraints simultaneously.
  3. 3.
    In PBEBench, programs modify on average only two examples, which is less reflective of real sound law induction problems, where many more words may be affected by each rule.
  4. 4.
    We evaluate only the strongest open and closed-source models on PBEBench. While this partly reflects limited time and computing resources, it also ensures the analysis is meaningful, as weaker models plateau much earlier and fail to match gpt-oss-120b and GPT-5 on longer cascades.
Appendix C Theoretical Framework
@@
We propose the function $\mathsf{feeds}(\cdot,\cdot)$, which classifies pairs of rules as feeding or not feeding.
@@
Definition C.3 ($\mathsf{Substr}$).
@@
Appendix D Benchmark Details
@@ D.2 Benchmark Statistics
Appendix E Method Details

### v4

@@ 5.3 PBEBench Performance
We evaluate the strongest open and closed-source models (gpt-oss-120b and GPT-5) on PBEBench across cascade lengths ranging from 2 to 30. GPT-5 consistently outperforms gpt-oss-120b, maintaining non-trivial Pass@1 at longer cascades where gpt-oss-120b performance collapses. At length 20, GPT-5 achieves moderate accuracy with a small sampling budget, while gpt-oss-120b drops to near-zero performance even with larger budgets. Performance declines monotonically with cascade length for both models, revealing clear reasoning limits. Trends across lengths are summarized in Fig. 10(a) and Fig. 10(b). A regression analysis further confirms cascade length as a dominant negative predictor of success, with the presence of specific BFCC relations associated with failure or success. Full results and analyses are deferred to Appendix F.1.
5.4 Ablations and Scaling
We perform ablations using PBEBench-Lite and its variant PBEBench-Lite-MoreEg (Appendix C.2) to analyze the effects of just increasing input-output examples on the gpt-oss models. Increasing the number of examples per PBE instance results in a slight reduction in performance. We also analyze the effect of the scaling strategies: sampling budget and thinking budget on GPT-5 and gpt-oss-120b on the harder PBEBench data. We observe consistent scaling behavior for both models, with rapid initial gains from increased compute, followed by diminishing returns and “hitting a wall” for longer ground truth cascades. Detailed quantitative results, scaling curves, and variance analyses are provided in Appendix F.2.
5.5 Real SLI Performance
We evaluate all open-source models on real SLI data (Appendix C.3), with results in Table 13. We additionally report performance on CLUTRR Sinha et al. (2019) and SLR-Bench Helff et al. (2025) in Appendix F.3. Motivated by our hypothesis that SLI is bottlenecked by multi-step inductive reasoning (Section 1), we measure Pearson rank correlations between model rankings on real SLI and those induced by PBEBench-Lite, CLUTRR, and SLR-Bench (Table 14). PBEBench-Lite shows the strongest correlation with real SLI ($r=0.8273$, $p=0.001677$), followed closely by SLR-Bench ($r=0.8091$, $p=0.002559$). However, SLR-Bench is sensitive to Prolog coding proficiency, with Codestral-22B achieving high syntax scores but weak overall performance relative to Qwen models, whose reasoning is stronger despite poorer Prolog syntax (Table 11). By relying on Python, PBEBench reduces syntax bottlenecks that would otherwise underestimate inductive capability.
@@ 6 Discussion
After analyzing the results, we address all the hypotheses in Section 1. H1: LLMs struggle with multi-step inductive reasoning. As discussed in Section 5.1, we see non-reasoning LLMs struggle to solve even simple multi-step PBE problems, and most open-source reasoning models outperform closed-source non-reasoning ones. Additionally, even the best performing reasoning models, GPT-5 and gpt-oss-120b, break when evaluated on the hard PBEBench data (Section 5.3) despite inference-time scaling strategies (Section 5.4). H2: Difficulty in our dataset is driven by interaction structure.: We find that increasing per-step difficulty by adding more examples has little effect on performance (Section 5.4). In contrast, factorial and logistic regression analyses (Sections 5.1,5.3) and performance degradation on longer cascades in PBEBench (Figs. 10(a),10(b)) show that LLMs struggle primarily with long cascades and complex BFCC interactions. Moreover, many models, including reasoning LLMs, fail to recover the correct program order in PBEBench-Lite-Perm even when given the correct shuffled programs for cascades of length 2 to 5, indicating that task difficulty arises from interactions within the cascade rather than individual steps. H3: Performance on PBEBench predicts performance on real SLI: As discussed in Section 5.5, even the simpler PBEBench-Lite has higher correlation with the model rankings on the real SLI data compared to CLUTRR and SLR-Bench, indicating the predictive power of our approach. H4: Scaling the thinking budget is more effective than increasing the sampling budget. A direct comparison for gpt-oss-120b shows that increasing sequence length with sufficient sampling is more time and cost-efficient than increasing sampling budget alone, as it reduces wasted attempts from unterminated chains of thought.
@@ 7 Conclusion and Future Work
We show that the proposed benchmark is challenging for many models, especially non-reasoning LLMs, because despite simple individual PBE steps, its sequential structure and BFCC-induced ordering constraints create emergent complexity that requires long chains of thought at test time. Increasing cascade length or interaction complexity breaks even the strongest reasoning models, including GPT-5 and gpt-oss-120b. It is also practically meaningful: PBEBench-Lite performance best predicts real SLI performance, suggesting that LLMs’ difficulty with multi-step inductive reasoning partly explains their poor real-world SLI results. In future work, we plan to synthesize progressively harder training data using PBEBench and explore hybrid neuro-symbolic data generation.
Limitations
Despite the contributions of this work, several limitations remain:
  1. 1.
    We primarily evaluate the strongest open and closed-source models on PBEBench. While constrained by time and compute, this choice stems from the analysis on PBEBench-Lite, where weaker models fail to match gpt-oss-120b or GPT-5, especially on longer cascades.
  2. 2.
    Owing to budget and time constraints, evaluations on CLUTRR, SLR-Bench, and real SLI data are limited to open-source models.
  3. 3.
    For PBEBench-Lite, PBEBench-Lite-MoreEg, CLUTRR, and SLR-Bench, we use a single sample per problem (sampling budget of 1), which may underestimate performance under larger sampling budgets.
  4. 4.
    For real SLI, we simplify the task by chunking cases with more than 200 examples into subsets of 50, relaxing program constraints by allowing longer cascades and larger substring rewrites. While this could enable degenerate solutions, most models still perform poorly even with increased thinking and sampling budgets.
  5. 5.
    Due to cost, GPT-5 is evaluated only on the most challenging cascade lengths and with limited sampling budgets, restricting comprehensive analysis across settings.
  6. 6.
    The random sampling procedure cannot perfectly balance constraints such as relation types and cascade lengths, particularly for long cascades where valid examples are rare, leading to high rejection rates and exhaustion of the patience parameter $\tau$. Future work could explore more efficient sampling strategies.
  7. 7.
    In PBEBench, individual programs modify relatively few examples on average, which is less representative of real sound law induction, where rules often apply to many words simultaneously.
Ethics Statement
Our work introduces a provably correct, fully automated, and scalable pipeline for generating multi-step Programming by Example (PBE) problems, enabling the evaluation of inductive reasoning in Large Language Models without human supervision. As no human subjects are involved (apart from the individuals who contributed the published data for the SLI task), we do not anticipate ethical concerns. Moreover, since the dataset focuses solely on assessing abstract inductive reasoning in LLMs, it is unlikely to be misused for harmful purposes.
Appendix
@@
    Appendix A discusses steps taken to ensure reproducibility.
@@
    Appendix B provides the theory and proofs for the relation type classifiers used to control the distribution.
@@
    Appendix C details the benchmark and metrics, including licensing, statistics, and related information.
@@
    Appendix D gives additional methodological details, including prompts, model selection, licensing, run costs, algorithm pseudocode, snapshots for closed-source models, and the K-fold evaluation used to simulate GPT-5’s sampling budget.
@@
    Appendix E outlines experimental details such as computational environment, inference parameters, and strategies for CoT truncation/control with gpt-oss-120b.
@@
    Appendix F presents additional results, including data generation efficiency, factorial and logistic regression analyses of instance difficulty, confusion matrices comparing dataset ground truth and model predictions, and extended tables for scaling experiments, effects of additional examples, and performance breakdowns by cascade length.
Appendix A Reproducibility Statement
To facilitate the reproduction of our results, we thoroughly document all components of our work. The BFCC relation types, their detectors (Algorithm 2), and proofs of correctness are presented in Appendix B. The rejection-sampling-based data generation process is both described (Section 3.1) and formally specified (Algorithm 1). Additionally, the algorithm for the feeding-bleeding swap permutation for creating the program reordering data is documented in Algorithm 3. Detailed statistics and parameters for all generated datasets appear in Section 4.1 and Appendix C.2. Our prompting and program extraction strategy, including scaling techniques, is reported in Section 3.2, while the prompt template is provided in Appendix D.3, and inference parameters for each model are given in Appendix E.2. For closed-source models, we use specific dated snapshots/checkpoints of models as recorded in Appendix D.7 to aid with reproducibility. To estimate the amount of variance between runs/experiments, we do a small controlled experiment for gpt-oss-120b (Table 26), and compute k-fold average for GPT-5 (Appendix D.8) The evaluation procedure and metrics are described in Section 4.3 and Section 3.2.
Appendix B Theoretical Framework
@@
We propose the function $\mathsf{feeds}(\cdot,\cdot)$ (equation 1), which classifies pairs of rules as feeding or not feeding.
@@
Definition B.3 ($\mathsf{Substr}$).
@@
Appendix C Benchmark Details
@@ C.2 Benchmark Statistics
C.3 Real SLI Benchmark
We construct a benchmark of 21 instances derived from real SLI data, covering 6 proto-language and attested-language pairs: prototangkhulic-huishu, austronesian-hawaiian, austronesian-niue, austronesian-tongan, austronesian-rarotongan, and austronesian-samoan. Our data is derived from Naik et al. (2025a, 2024) original datasets. Each instance consists of 50 example input and output pairs and uses the same prompt template as the multi-step PBE task prompt employed for all synthetic datasets described above. However, the prompt constraints permit program cascades of up to 50 programs, the sizes of $\alpha_{k}$ and $\beta_{k}$ can be up to 5, and each LLM is allotted 32 solution attempts (sampling budget) per PBE instance to better reflect the real-world complexity of low-resource SLI. We do note a limitation of these assumptions, as for some language pairs, such as prototangkhulic-huishu, certain input words are shorter than 5 characters, and in principle an LLM could exploit this by constructing a single string rewrite rule that replaces the entire input word with the target output word. Nevertheless, despite this potential for gaming, the results indicate that most LLMs struggle with the task, and even the best-performing gpt-oss-120B model achieves a Pass@1 of only $0.33$, despite a sampling budget of 32.
C.4 Metrics Details
For the multi-step PBE task, since a given list of inputs ($\vec{\imath}$) could be transformed into the outputs ($\vec{o}$) by multiple program cascades $\vec{p}$, we utilize metrics based on functional equivalence. This is in line with programming by example literature Li and Ellis (2024) and historical linguistics Hoenigswald (1960).
Concretely, we execute the model-generated solution $\hat{\vec{p}}$ on the inputs, treating the input-output pairs as test cases. We then compare the predicted outputs $\hat{\vec{o}}=\hat{\vec{p}}(\vec{\imath})$ with the ground-truth outputs $\vec{o}$ at two different levels of granularity: (1) coarse-grained evaluation, corresponding to solve rate (Pass@1), and (2) fine-grained evaluation, corresponding to normalized edit similarity (Edit_Sim).
Both metrics perform element-wise comparisons on the strings in the output vectors:
Coarse-grained Metric (Pass@1): This metric corresponds to Pass@1 as introduced in Chen et al. (2021b).
$\mathrm{pass@1}=\frac{1}{|\mathcal{D}|}\sum_{\vec{o},\vec{\imath}\in\mathcal{D}}\mathrm{1}_{\hat{\vec{p}}(\vec{\imath})=\vec{o}}$
Here $\mathrm{1}_{\hat{\vec{p}}(\vec{\imath})=\vec{o}}$ is an indicator variable which equals 1 when $\hat{\vec{p}}(\vec{\imath})=\vec{o}$ and 0 otherwise.
Fine-grained Metric (Edit_Sim): This metric is identical to the Reward@1 metric used by Naik et al. (2025b).
$\mathrm{Edit\_Sim}=\frac{1}{|\mathcal{D}|}\sum_{\vec{o},\vec{\imath}\in\mathcal{D}}1-\frac{\mathrm{dist}(\hat{\vec{p}}(\vec{\imath}),\vec{o})}{\mathrm{dist}(\vec{\imath},\vec{o})}$
Here $\mathrm{dist}$ denotes the total Levenshtein edit distance, summed across the corresponding strings in the predicted and ground-truth output vectors.
In addition to these performance metrics, we evaluate the proportion of programs generated by an LLM that are valid, meaning they satisfy all constraints specified in the prompt (see Section 3.2). We refer to this metric as the Valid_Rate.
Finally, for the program reordering task, given a model-predicted permutation $\hat{\sigma}$, we apply it to the originally permuted program sequence $\sigma(\vec{p})$ and evaluate whether it correctly recovers the original outputs. Specifically, we execute $\hat{\sigma}(\sigma(\vec{p}))$ on the inputs and check whether $\hat{\sigma}(\sigma(\vec{p}))(\vec{\imath})=\vec{o}$.
For some instances, we explicitly construct the problem such that there exists only a single unique solution, in which case the inverse permutation $\hat{\sigma}=\sigma^{-1}$ is the only correct answer. The evaluation metric used is accuracy (Acc), defined analogously to Pass@1:
$\mathrm{Acc}=\frac{1}{|\mathcal{D}|}\sum_{\vec{o},\vec{\imath}\in\mathcal{D}}\mathrm{1}_{\hat{\sigma}(\sigma(\vec{p}))(\vec{\imath})=\vec{o}}$
We additionally report unique accuracy, defined as the accuracy computed over the subset of instances with a single valid solution. In the PBEBench-Lite-Perm dataset, this subset consists of 242 instances.
@@ C.4 Metrics Details
Appendix D Method Details
@@ D.1 Algorithm Pseudocode
D.2 Program Extraction
To extract the program cascade from LLM predictions, we apply the following rules. If the LLM fails to produce a valid, extractable Python code block, or produces a null response due to unterminated chain of thought (as observed in gpt-oss models), we mark the corresponding case as a null response and replace the prediction with an identity program like replace("x", "x") for evaluation. For cases where a valid response is produced, we account for the fact that we evaluate a wide range of LLMs, including reasoning models that may generate intermediate programs and subsequently refine or improve them through reflection. Accordingly, we evaluate both the first and the last Python code blocks produced by the model. For the remaining valid cases, if a predicted program cascade $\hat{\vec{p}}$ contains more than $L_{max}$ programs, we retain only the first $L_{max}$ programs for evaluation. If the predicted program violates any other constraint, it is replaced by an identity program $p^{I}$, which leaves the inputs unchanged. Given the predicted cascade $\hat{\vec{p}}$, we compute the predicted outputs as $\hat{\vec{o}}=\hat{\vec{p}}(\vec{\imath})$. Table 3 reports performance on PBEBench-Lite for both the first and last code blocks. We observe that, for most models, the last code block yields superior performance. Therefore, in Table 2 in the main paper, as well as in any table where it is not explicitly specified otherwise, we report results corresponding to the last code block by default.
Model First Code Block Last Code Block
Pass@1 Edit Sim Valid Rate Pass@1 Edit Sim Valid Rate
Codestral-22B 0.0109 -0.0130 0.8280 0.0109 -0.0107 0.8254
Qwen2.5-32B-Instruct 0.0298 0.1232 0.8265 0.0298 0.1252 0.8288
Qwen3-32B 0.0179 0.0964 0.7645 0.0179 0.0964 0.7645
Qwen3-30B-A3B ★◼ 0.2887 0.3360 0.9905 0.2887 0.3360 0.9905
QwQ-32B ★ 0.3591 0.4072 0.9500 0.3601 0.4092 0.9493
Qwen3-32B (with CoT) ★ 0.3938 0.4803 0.9182 0.4187 0.5026 0.9676
gpt-oss-20b ★◼ 0.4058 0.4619 0.9900 0.4058 0.4619 0.9900
gpt-oss-120b ★◼ 0.6250 0.6985 0.9254 0.6250 0.6985 0.9254
Claude-3.5-Sonnet 0.1845 0.4430 0.8208 0.1845 0.4434 0.8209
Claude-3.7-Sonnet 0.2212 0.4825 0.8409 0.2321 0.4996 0.8409
Claude-4 Sonnet 0.2956 0.5832 0.7720 0.2966 0.5870 0.7719
o3-mini ★ 0.5278 0.5954 0.9179 0.5278 0.5954 0.9179
o4-mini ★ 0.6329 0.6907 0.9153 0.6329 0.6907 0.9153
Gemini 2.5 Flash ★ 0.5833 0.6533 0.7911 0.5863 0.6562 0.7901
GPT-5 ★ 0.7242 0.7645 0.9286 0.7242 0.7645 0.9286